OpenStudy (anonymous):

please help. The amount a car dealership earns for selling a car is directly proportional to the price of the car. The table shows this relationship. What is the constant of proportionality? Car price ($) (x) 20,000 30,000 40,000 Amount earned ($) (y) 1000 1500 2000 Answer = ? (medal)

OpenStudy (anonymous):

@muscrat123 can you help

OpenStudy (lulubj):

0.05

OpenStudy (anonymous):

how did you get that

OpenStudy (muscrat123):

yes how did u get that

OpenStudy (anonymous):

ok thanks

OpenStudy (muscrat123):

do u know how to get the answer?

OpenStudy (anonymous):

not really

OpenStudy (muscrat123):

simply divide x by y

OpenStudy (muscrat123):

:)

OpenStudy (lulubj):

nott 20 nono

OpenStudy (lulubj):

its 1/20 so 0.05

OpenStudy (lulubj):

x = 20,000 y = 1,000 y/x = 1,000/20,000 = 0.05 y = 0.05 x The constant of proportionality is 0.05 (5 percent). Look at the other points just to check it out and see if it holds water: x = 1,500 y = 30,000 x/y = 0.05 yes ! x = 2,000 y = 40,000 x/y = 0.05 yes !

OpenStudy (lulubj):

@muscrat123 you gave him the wrong answer!

OpenStudy (muscrat123):

no, u divided wrong

OpenStudy (muscrat123):

the wrong #s

OpenStudy (muscrat123):

its flipped

OpenStudy (lulubj):

no, i didnt

OpenStudy (muscrat123):

it is x / y

OpenStudy (anonymous):

its ok guys i all ready entered it in and i failed

OpenStudy (lulubj):

told you! look at the other points and how i also got 0.05